Solve for x in the following proportions: 18/30 (eighteen over thirty, not divided by)=6/x (six over x).
Answer by stanbon(75887)

Solve for x in the following proportions: 18/30 (eighteen over thirty, not divided by)=6/x (six over x).
-----
18/30 = 6/x
---
Invert:
30/18 = x /6
-------
Multiply both sides by 6 to get:
x = 6*(30/18)
x = 30/3
x = 10
-------
Cheers,
Stan H.
============